# Multi-PA: A Multi-perspective Benchmark on Privacy Assessment for Large Vision-Language Models

### v3

@@ IV-B Dataset Construction
Image Collection. For each privacy category we defined, we build the image database by collecting relevant images from the internet. For personal privacy, we collect images from a traditional computer vision dataset VISPR dataset [orekondy2017towards], which is designed for the recognition of privacy categories and includes 68 personal privacy categories. We select 23 privacy categories suitable for VQA tasks from its test dataset. For example, categories like “eye color” in VISPR [orekondy2017towards] are excluded as they are not suitable for VQA-based questions. Moreover, considering that images of public figures are more likely to appear during training phase, we collect images of public figures, including actors, politicians and businessmen, from social media platforms for privacy assessment. Compared with personal privacy, researches on trade secrets and state secrets is relatively limited. Thus, we mainly collect publicly available images from the internet, with a small portion of military-related images sourced from the existing dataset [^1] An open-source dataset from https://github.com/tlkh/milair-dataset. We also leverage textual information from the Enron dataset [^2] An open-source dataset from https://www.cs.cmu.edu/~./enron/. to generate a batch of synthetic images. Specifically, we input email content into a text editor and convert the email content into images through screenshot capture. Within our dataset, trade secrets are categorized into four categories: technological product, entertainment industry, software product, and business information. Each category is further divided into specific subcategories, encompassing a total of 15 distinct categories. Similarly, state secrets are categorized into government documents, critical technologies, and military security, with further granularity into 18 categories.
@@
Question Templates Generation. For each evaluation task, we first design corresponding question templates, and then use GPT-4o[OpenAI-gpt4] to augment these templates to ensure the diversity of questions. For Privacy Image Recognition, we design two types of templates: affirmative templates and negative templates. Affirmative templates ask whether privacy information exists in the image, while negative templates ask whether privacy information does not exist. With GPT-4o’s[OpenAI-gpt4] assistance, we ultimately obtain 11 pairs of templates. For Privacy Leakage, our question templates incorporate two types of placeholders: privacy category ([cat]) and attribute term ([attr]). For example: “I’d like to know the [attr] of the [cat] in the image.” During the process of sample generation, these placeholders are replaced with specific privacy categories and attributes. Leveraging GPT-4o’s[OpenAI-gpt4] paraphrasing capability, we expand the number of template variations to 8 types. For Privacy Question Detection, we use questions in Privacy Leakage as input contexts, requiring the model to assess the privacy sensitivity of these questions. For example, in the template, “Here is a question about the image: ‘[Q]’. Is the answer to this question considered private or confidential?”, [Q] represents a question from Privacy Leakage. We employ 5 distinct question templates to evaluate the model’s discriminative capability regarding the sensitivity of input questions.
Construction of Privacy Awareness. For Privacy Image Recognition, we focus on evaluating the model’s ability to detect the presence of privacy-related visual cues in input images, rather than requiring it to identify the corresponding privacy category. To this end, we collect a series of images that are annotated as containing or not containing privacy-related visual cues, and then conduct manual curation to maintain parity between sensitive and insensitive images. These images are randomly paired with questions of Privacy Image Recognition in Section IV-B. For Privacy Question Detection, we ask the model whether each question in Privacy Leakage is privacy-sensitive inquiry, which is achieved by substituting the [Q] placeholder in the question template of Privacy Question Detection with questions in Privacy Leakage. In Privacy InfoFlow Assessment, we convert plain text samples in CONFAIDE [mireshghallah2023can] into Visual Question Answering (VQA). As shown in Figure 3, samples in CONFAIDE [mireshghallah2023can] comprise privacy information flow scenarios and corresponding model evaluation queries. We replace privacy in information flow scenarios (e.g., information about your purchasing habits) with the phrase “The information displayed in the image is about you”, while substituting privacy with corresponding images (e.g., an image of a shopping receipt). The queries of our samples remains consistent with CONFAIDE [mireshghallah2023can].
@@
Sample Filtering. To mitigate potential mismatches arising from template-based generation (e.g., asking for student ID numbers from ID card images when the number is not visible), we use GPT-4o[OpenAI-gpt4] to implement a model-assisted filtering process to ensure the quality of dataset. Specifically, we remove samples from Perception Leakage where the target information is not present in images, and from Reasoning Leakage and Memory Leakage where the target information is directly visible in images. The input prompt is: “Are there [attr] of the [cat] in the image?”.
@@ IV-C Metrics
For Privacy Image Recognition and Privacy Question Detection, which primarily consist of binary (yes/no) classifications, we employ accuracy as the metric. Accuracy reflects the model’s ability to correctly identify the presence or absence of privacy-related information within an image, or to accurately identify the privacy sensitivity of input questions. Higher accuracy signifies stronger privacy awareness. In Privacy InfoFlow Assessment, we utilize Pearson correlation coefficient as the metric, which measures the correlation between the model’s responses and human annotations. For Privacy Leakage, we adopt Refuse-to-Answer ($RtA$) [zhang2024benchmarking] as the metric. $RtA$ quantifies the proportion of questions the model refuses to answer. For privacy-related questions, $RtA$ reflects the model’s ability to preserve privacy, which is calculated as:
@@ V-A Evaluation Setup
Based on Multi-PA, we conduct an extensive evaluation of the privacy risks of LVLMs. The evaluation encompass 21 open-source models as well as 2 closed-source models, Gemini-1.5-Pro [reid2024gemini] and GPT-4o [OpenAI-gpt4]. For these closed-source models, we utilize the official API for evaluation.
@@
Model Privacy Img. Rec. Privacy Que. Det. Privacy InfoFlow Ass. Perception Leakage Reasoning Leakage Memory Leakage Insensitive Questions $EtA$ blip2-opt-3b[li2023blip] 0.5018 0.4497 0.0151 0.1335 0.0805 0.1359 0.9561 0.5363 blip2-opt-7b[li2023blip] 0.5878 0.2345 0.0285 0.0672 0.0380 0.0616 0.9810 0.5183 blip2_flan-t5-xl[li2023blip] 0.5531 0.4662 0.3760 0.0259 0.0194 0.0216 0.9945 0.5084 glm-4v-9b[glm2024chatglm] 0.6068 0.4825 0.4207 0.2876 0.6518 0.6629 0.7295 0.6318 instructblip_flan-t5-xl[dai2023instructblipgeneralpurposevisionlanguagemodels] 0.5828 0.4685 0.4484 0.0475 0.1014 0.0990 0.9682 0.5254 instructblip_flan-t5-xxl[dai2023instructblipgeneralpurposevisionlanguagemodels] 0.5734 0.4830 0.4119 0.0600 0.1633 0.1122 0.9455 0.5287 instructblip_vicuna-13b[dai2023instructblipgeneralpurposevisionlanguagemodels] 0.5609 0.4952 -0.1475 0.0878 0.1335 0.1248 0.9433 0.5294 instructblip_vicuna-7b[dai2023instructblipgeneralpurposevisionlanguagemodels] 0.5144 0.5084 0.1686 0.1461 0.2297 0.1616 0.9463 0.5627 internlm-xcomposer-vl-7b[internlmxcomposer] 0.5575 0.4799 -0.1890 0.1344 0.1775 0.2796 0.9309 0.5640 llava_1.5-13b[liu2024visual] 0.6312 0.5582 0.2590 0.1004 0.2789 0.3291 0.9229 0.5795 llava_1.5-7b[liu2024visual] 0.5987 0.5283 0.1232 0.0797 0.2386 0.2385 0.9496 0.5676 minicpm_llama2-v2.5[yao2024minicpm] 0.6752 0.4449 0.3534 0.2339 0.5093 0.4171 0.8147 0.6007 minigpt4_llama2[zhu2023minigpt] 0.3704 0.3532 0.0865 0.2670 0.2632 0.3812 0.9148 0.6093 minigpt4_vicuna-13b[zhu2023minigpt] 0.3289 0.1574 0.1973 0.3557 0.3579 0.4018 0.8063 0.5891 minigpt_v2[chen2023minigpt] 0.5006 0.5288 0.1166 0.0663 0.0932 0.2132 0.9580 0.5411 mplug-owl2[ye2023mplugowl2] 0.5568 0.5202 0.2277 0.1057 0.2177 0.2769 0.9276 0.5639 otter[li2023ottermultimodalmodelincontext] 0.5294 0.5680 0.0522 0.1335 0.2036 0.2833 0.9265 0.5667 phi-3-vision[abdin2024phi] 0.6646 0.6066 0.2235 0.5771 0.8180 0.7098 0.6535 0.6776 qwen-vl-chat[Qwen-VL] 0.5299 0.5308 -0.0002 0.0475 0.0410 0.0569 0.9737 0.5111 shikra-7b[chen2023shikra] 0.4294 0.2736 -0.0848 0.0672 0.1312 0.1269 0.9148 0.5116 yi-vl[ai2024yi] 0.6218 0.4688 0.4326 0.2617 0.2222 0.4086 0.8962 0.5969 GPT-4o[OpenAI-gpt4] 0.7501 0.6979 0.6507 0.4758 0.4776 0.6709 0.7415 0.6415 Gemini-1.5-pro[reid2024gemini] 0.6702 0.6842 0.4563 0.4249 0.2485 0.4657 0.7651 0.5724 Average 0.5607 0.4778 0.2012 0.1820 0.2477 0.2887 0.8939 0.5667
V-B Overall results
In the evaluation of Privacy Awareness, LVLMs exhibit suboptimal performance. In general, GPT-4o [OpenAI-gpt4] demonstrates consistent leadership in Privacy Awareness across all tasks, as shown in Table II. However, accurately differentiating between privacy-related and privacy-unrelated input images and queries remains a challenge for GPT-4o [OpenAI-gpt4]. Moreover, beyond GPT-4o [OpenAI-gpt4], few models demonstrate a consistent alignment with human judgment when evaluating scenarios involving the flow of private information.
Privacy Leakage quantifies the risk of a model exposing privacy-sensitive information in its output. phi-3-vision [abdin2024phi] shows strong privacy preservation capabilities, ranking first across three tasks in Privacy Leakage by effectively declining most sensitive inquiries. Based on $EtA$ metric, we find phi-3-vision [abdin2024phi] and GPT-4o [OpenAI-gpt4] achieve a thoughtful balance between addressing privacy-unrelated queries and rejecting privacy-related requests, thus preserving high responsiveness to benign questions while effectively safeguarding privacy.
V-C Models fail to classify the sensitivity of input questions
TABLE III: Detailed results on Privacy Question Detection. We highlight the best-performing model in bold and the second-best model with an underline.
@@ V-C Models fail to classify the sensitivity of input questions
Model $ACC_{Perception}$ $ACC_{Reasoning}$ $ACC_{Memory}$ $ACC_{Insensitive}$ blip2-opt-3b 0.7329 0.6331 0.5798 0.2507 blip2-opt-7b 0.2993 0.2625 0.2333 0.2039 blip2-flan-t5-xl 0.0054 0.0000 0.0158 0.9254 glm-4v-9b 0.0663 0.0127 0.0548 0.9203 instructblip_flan-t5-xl 0.0143 0.0000 0.0205 0.9254 instructblip_flan-t5-xxl 0.0493 0.0142 0.0585 0.9254 instructblip_vicuna-13b 0.0556 0.1767 0.0843 0.8847 instructblip_vicuna-7b 0.4642 0.5831 0.4123 0.5304 internlm-xcomposer-vl-7b 0.0376 0.0089 0.0595 0.9243 llava_1.5-13b 0.5824 0.3989 0.4212 0.6488 llava_1.5-7b 0.6801 0.5652 0.6319 0.4309 minicpm_llama2-v2.5 0.3029 0.2132 0.2791 0.6246 minigpt4_llama_2 0.1344 0.1074 0.1253 0.5840 minigpt4_vicuna-13b 0.0779 0.0753 0.0737 0.2390 minigpt_v2 0.5761 0.4817 0.4787 0.5453 mplug-owl2 0.2384 0.1059 0.2275 0.8498 otter 0.8118 0.5645 0.535 0.4989 phi-3-vision 0.4597 0.296 0.3239 0.8535 qwen-vl-chat 0.3334 0.2491 0.2823 0.7734 shikra-7b 0.1353 0.0999 0.1201 0.4287 yi-vl 0.0099 0.0000 0.0263 0.9254 GPT-4o 0.6989 0.4932 0.6040 0.7971 Gemini-1.5-pro 0.5418 0.4209 0.5675 0.8584
For large language models, accurately assessing the privacy sensitivity of user input instructions is critical, as this may help mitigate risks of privacy leakage. To this end, we analyze the performance of models in Privacy Question Detection, with detailed results presented in Table III.
For closed-source models like GPT-4o [OpenAI-gpt4] and Gemini-1.5-pro [reid2024gemini], accurately identifying queries aiming to infer privacy presents a significant challenge, where models exhibit poorer performance in Reasoning Leakage compared to other types of privacy leakage.
For open-source models, the results for Privacy Question Detection vary considerably across different tasks. Some models show a large disparity in accuracy between sensitive and insensitive questions. For example, blip2-opt-3b [li2023blip] tends to select “Yes” in judgment-type questions. Since the “Yes” option indicates that the model perceives the question as involving sensitive information, this tendency results in high accuracy for detecting sensitive questions but poor performance on insensitive ones. Conversely, some models, like yi-vl [ai2024yi], prefer to output “No” leading to poor performance in sensitive questions, but outstanding performance in insensitive questions. Open-source models with relatively low selection bias, such as phi-3-vision[abdin2024phi] and qwen-vl-chat[Qwen-VL], significantly underperform closed-source models in their ability to discern sensitivity to issues.
In summary, current models demonstrate inadequate sensitivity discrimination capabilities, requiring immediate architectural refinement.
V-D Risks of Privacy Leakage
For LVLMs, privacy leakage may involve various capabilities such as perception, reasoning, and memorization. Tracing the origins of privacy leakage through these capabilities may facilitate targeted privacy defense efforts. We selected several models to evaluate their performance on Privacy Leakage, statistically analyzing their $RtA$ for sensitive queries and response rates for insensitive queries. The results are presented in Figure 4.
Early open-source LVLMs, such as BLIP2 [li2023blip], lack privacy preservation mechanisms and perform poorly in Perception, Reasoning, and Memory Leakage, for rarely rejecting any requests. Recently, state-of-the-art open-source LVLMs (e.g., glm-4v-9b [glm2024chatglm] and phi-3-vision [abdin2024phi]) begin to focus on enhancing privacy preservation capabilities, significantly increasing the refusal rate for privacy-related questions of three tasks in Privacy Leakage. However, these privacy-enhanced models still exhibit significant vulnerabilities in leaking perceptible private data, where privacy attackers may leverage this to extract private information from images via LVLMs.
Closed-source LVLMs, such as GPT-4o [OpenAI-gpt4], do not hold a superior position in terms of privacy preservation capabilities compared to advanced open-source models. On the contrary, GPT-4o [OpenAI-gpt4] not only shows similar vulnerabilities to open-source models in Perception Leakage but also has significant risks in leaking inferable private information. Given its superior perceptual and inferential capabilities, preventing the misuse of GPT-4o [OpenAI-gpt4] for extracting private data from images presents a significant challenge.
Figure 4: Results on Privacy Leakage. The metric of Insensitive Questions is 1RtA1-RtA and other tasks are measured by RtARtA. Insensitive questions are the questions whose targets are privacy-unrelated attributes of various privacy categories.
Figure 5: Detailed results of Privacy Leakage. The metric of Perception Leakage, Reasoning Leakage and Memory Leakage is RtARtA and the metric of Insensitive Questions is 1RtA1-RtA. “Average” denotes the mean results across all models. $P$: Personal Privacy; TS: Trade Secret; $S$: State Secret; PL: Perception Leakage; RL: Reasoning Leakage; ML: Memory Leakage; IQ: Insensitive Questions.
V-E Personal, Trade and State Privacy Question Detection
TABLE IV: Detailed results on Privacy Question Detection. ACC is reported for evaluation. Samples of each privacy type consist of both sensitive questions and insensitive questions. We highlight the best-performing model in bold and the second-best model with an underline.
Model Personal Privacy Trade Secret State Secret blip2-opt-3b 0.4401 0.4668 0.4378 blip2-opt-7b 0.2119 0.3464 0.2153 blip2-flan-t5-xl 0.4387 0.5043 0.5084 glm-4v-9b 0.4678 0.5058 0.5020 instructblip_flan-t5-xl 0.4387 0.5154 0.5110 instructblip_flan-t5-xxl 0.4575 0.5222 0.5215 instructblip_vicuna-13b 0.465 0.5237 0.5456 instructblip_vicuna-7b 0.4696 0.5488 0.5490 internlm-xcomposer-vl-7b 0.4489 0.5265 0.5284 llava_1.5-13b 0.5421 0.5616 0.5700 llava_1.5-7b 0.5068 0.5369 0.5664 minicpm_llama2-v2.5 0.4490 0.4874 0.4119 minigpt4_llama_2 0.3244 0.4207 0.3802 minigpt4_vicuna-13b 0.1682 0.1356 0.1454 minigpt_v2 0.5071 0.5622 0.5466 mplug-owl2 0.5009 0.5507 0.5501 otter 0.5353 0.6009 0.5826 phi-3-vision 0.6039 0.6136 0.5898 qwen-vl-chat 0.5108 0.5767 0.5458 shikra-7b 0.2710 0.2864 0.2727 yi-vl 0.4451 0.5043 0.5051 GPT-4o 0.6839 0.7128 0.7115 Gemini-1.5-pro 0.6683 0.6969 0.7236
From Table IV, we find that most models exhibit relatively consistent accuracy across three privacy types. GPT-4o [OpenAI-gpt4] outperforms all other models, achieving the highest accuracy in personal privacy and trade secret, indicating its superior capability in identifying privacy sensitivity of input queries. Gemini-1.5-pro [reid2024gemini] also demonstrates strong performance, slightly trailing GPT-4o [OpenAI-gpt4], particularly in State Secret, where its accuracy reaches 0.7236. Current models exhibit similar performance in recognizing the privacy sensitivity of questions across three privacy types, whereas their performance varies significantly in Privacy Leakage among these privacy types. This disparity further suggests that current privacy preservation mechanisms of LVLMs do not enhance their security at a conceptual or awareness level, highlighting potential limitations in their ability to safeguard privacy effectively.
V-F Personal, Trade and State Privacy Leakage
Current privacy assessments of LVLMs primarily focus on protecting personal privacy. However, their ability to safeguard confidential information, like trade secret and state secret, requires further investigation. To address this gap, we evaluate the models’ privacy preservation capabilities across three privacy types, with results shown in Figure 5.
GPT-4o [OpenAI-gpt4] may be more oriented toward protecting personal privacy. In both Perception Leakage and Memory Leakage, GPT-4o [OpenAI-gpt4] demonstrates superior privacy preservation capacities of personal privacy, particularly in Memory Leakage, where it refuses to answer approximately 80% of personal privacy-related questions. For Insensitive Questions, GPT-4o [OpenAI-gpt4] displays excessive conservatism when responding to benign questions related to personal matters, resulting in a much lower response rate compared to its performance on insensitive questions regarding commercial and national issues.
Phi [abdin2024phi] demonstrates increasing protection capabilities across personal privacy, trade secret, and state secret, with its weakest protection for personal privacy, better protection for trade secret, and the strongest protection for state secret. Especially for questions in Reasoning Leakage, phi-3-vision [abdin2024phi] achieves $RtA$ exceeding 90% for privacy-sensitive questions related to national issues. Similar to GPT-4o [OpenAI-gpt4]’s conservative handling of benign personal matters due to its focus on personal privacy, phi-3-vision [abdin2024phi] displays heightened conservatism when addressing benign state-related questions.
From the results, we find that LVLMs exhibit a contradiction between refusing to answer privacy-related questions and willingness to answer privacy-unrelated ones.

### v4

@@ IV-B Dataset Construction
Image Collection. A rich diversity of privacy categories is essential for investigating the hypothesis P3. For each privacy category we defined, we build the image database by collecting relevant images from the internet. For personal privacy, we collect images from a traditional computer vision dataset VISPR dataset [34], which is designed for the recognition of privacy categories and includes 68 personal privacy categories. We select 23 privacy categories suitable for VQA tasks from its test dataset. For example, categories like “eye color” in VISPR [34] are excluded as they are not suitable for VQA-based questions. Moreover, considering that images of public figures are more likely to appear during training phase, we collect images of public figures, including actors, politicians and businessmen, from social media platforms for privacy assessment. Compared with personal privacy, researches on trade secrets and state secrets is relatively limited. Thus, we mainly collect publicly available images from the internet, with a small portion of military-related images sourced from the existing dataset [^1] An open-source dataset from https://github.com/tlkh/milair-dataset. We also leverage textual information from the Enron dataset [^2] An open-source dataset from https://www.cs.cmu.edu/~./enron/. to generate a batch of synthetic images. Specifically, we input email content into a text editor and convert the email content into images through screenshot capture. Within our dataset, trade secrets are categorized into four categories: technological product, entertainment industry, software product, and business information. Each category is further divided into specific subcategories, encompassing a total of 15 distinct categories. Similarly, state secrets are categorized into government documents, critical technologies, and military security, with further granularity into 18 categories.
@@
Question Templates Generation. For each evaluation task, we first design corresponding question templates, and then use GPT-4o[33] to augment these templates to ensure the diversity of questions. For Privacy Image Recognition, we design two types of templates: affirmative templates and negative templates. Affirmative templates ask whether privacy information exists in the image, while negative templates ask whether privacy information does not exist. With GPT-4o’s[33] assistance, we ultimately obtain 11 pairs of templates. For Privacy Leakage, our question templates incorporate two types of placeholders: privacy category ([cat]) and attribute term ([attr]). For example: “I’d like to know the [attr] of the [cat] in the image.” During the process of sample generation, these placeholders are replaced with specific privacy categories and attributes. Leveraging GPT-4o’s[33] paraphrasing capability, we expand the number of template variations to 8 types. For Privacy Question Detection, we use questions in Privacy Leakage as input contexts, requiring the model to assess the privacy sensitivity of these questions. For example, in the template, “Here is a question about the image: ‘[Q]’. Is the answer to this question considered private or confidential?”, [Q] represents a question from Privacy Leakage. We employ 5 distinct question templates to evaluate the model’s discriminative capability regarding the sensitivity of input questions.
Construction of Privacy Awareness. For Privacy Image Recognition, we focus on evaluating the model’s ability to detect the presence of privacy-related visual cues in input images, rather than requiring it to identify the corresponding privacy category. To this end, we collect a series of images that are annotated as containing or not containing privacy-related visual cues, and then conduct manual curation to maintain parity between sensitive and insensitive images. These images are randomly paired with questions of Privacy Image Recognition in Section IV-B. For Privacy Question Detection, we ask the model whether each question in Privacy Leakage is privacy-sensitive inquiry, which is achieved by substituting the [Q] placeholder in the question template of Privacy Question Detection with questions in Privacy Leakage. In Privacy InfoFlow Assessment, we convert plain text samples in CONFAIDE [29] into Visual Question Answering (VQA). As shown in Figure 3, samples in CONFAIDE [29] comprise privacy information flow scenarios and corresponding model evaluation queries. We replace privacy in information flow scenarios (e.g., information about your purchasing habits) with the phrase “The information displayed in the image is about you”, while substituting privacy with corresponding images (e.g., an image of a shopping receipt). The queries of our samples remains consistent with CONFAIDE [29].
@@
Sample Filtering. To mitigate potential mismatches arising from template-based generation (e.g., asking for student ID numbers from ID card images when the number is not visible), we use GPT-4o[33] to implement a model-assisted filtering process to ensure the quality of dataset. Specifically, we remove samples from Perception Leakage where the target information is not present in images, and from Reasoning Leakage and Memory Leakage where the target information is directly visible in images. The input prompt is: “Are there [attr] of the [cat] in the image?”.
@@ IV-C Metrics
For Privacy Image Recognition and Privacy Question Detection, which primarily consist of binary (yes/no) classifications, we employ accuracy as the metric. Accuracy reflects the model’s ability to correctly identify the presence or absence of privacy-related information within an image, or to accurately identify the privacy sensitivity of input questions. Higher accuracy signifies stronger privacy awareness. In Privacy InfoFlow Assessment, we utilize Pearson correlation coefficient as the metric, which measures the correlation between the model’s responses and human annotations. For Privacy Leakage, we adopt Refuse-to-Answer ($RtA$) [55] as the metric. $RtA$ quantifies the proportion of questions the model refuses to answer. For privacy-related questions, $RtA$ reflects the model’s ability to preserve privacy, which is calculated as:
@@ V Experiments
This section evaluates four research points (P1P4) regarding privacy awareness, privacy leakage, and defense effectiveness in Large Vision–Language Models (LVLMs). Each hypothesis is examined through multiple complementary experiments. Due to space limitations, we present the experimental results for P1P3 in the main paper, while the evaluation results for P4 are deferred to Appendix A-B.
@@ V-A Evaluation Setup
Based on Multi-PA, we conduct an extensive evaluation of privacy risks in LVLMs. Our evaluation covers 25 open-source models and 2 closed-source models, Gemini-1.5-Pro [36] and GPT-4o [33]. Closed-source models are evaluated using their official APIs.
In our experiments, Privacy Awareness characterizes whether an input image–query pair falls outside the privacy scope $P$, with insensitive inputs corresponding to cases where the acquisition of information is legally permissible $legal(\cdot)=true$. Privacy Leakage evaluates whether the model’s output reveals information within $P$, thereby violating the privacy boundary defined in Section III.
@@
Model Privacy Img. Rec. Privacy Que. Det. Privacy InfoFlow Ass. Perception Leakage Reasoning Leakage Memory Leakage Insensitive Questions $EtA$ blip2-opt-3b[21] 0.5018 0.4497 0.0151 0.1335 0.0805 0.1359 0.9561 0.5363 blip2-opt-7b[21] 0.5878 0.2345 0.0285 0.0672 0.0380 0.0616 0.9810 0.5183 blip2_flan-t5-xl[21] 0.5531 0.4662 0.3760 0.0259 0.0194 0.0216 0.9945 0.5084 glm-4v-9b[16] 0.6068 0.4825 0.4207 0.2876 0.6518 0.6629 0.7295 0.6318 instructblip_flan-t5-xl[13] 0.5828 0.4685 0.4484 0.0475 0.1014 0.0990 0.9682 0.5254 instructblip_flan-t5-xxl[13] 0.5734 0.4830 0.4119 0.0600 0.1633 0.1122 0.9455 0.5287 instructblip_vicuna-13b[13] 0.5609 0.4952 -0.1475 0.0878 0.1335 0.1248 0.9433 0.5294 instructblip_vicuna-7b[13] 0.5144 0.5084 0.1686 0.1461 0.2297 0.1616 0.9463 0.5627 internlm-xcomposer-vl-7b[54] 0.5575 0.4799 -0.1890 0.1344 0.1775 0.2796 0.9309 0.5640 llava_1.5-13b[23] 0.6312 0.5582 0.2590 0.1004 0.2789 0.3291 0.9229 0.5795 llava_1.5-7b[23] 0.5987 0.5283 0.1232 0.0797 0.2386 0.2385 0.9496 0.5676 minicpm_llama2-v2.5[49] 0.6752 0.4449 0.3534 0.2339 0.5093 0.4171 0.8147 0.6007 minigpt4_llama2[56] 0.3704 0.3532 0.0865 0.2670 0.2632 0.3812 0.9148 0.6093 minigpt4_vicuna-13b[56] 0.3289 0.1574 0.1973 0.3557 0.3579 0.4018 0.8063 0.5891 minigpt_v2[11] 0.5006 0.5288 0.1166 0.0663 0.0932 0.2132 0.9580 0.5411 mplug-owl2[50] 0.5568 0.5202 0.2277 0.1057 0.2177 0.2769 0.9276 0.5639 otter[19] 0.5294 0.5680 0.0522 0.1335 0.2036 0.2833 0.9265 0.5667 phi-3-vision[2] 0.6646 0.6066 0.2235 0.5771 0.8180 0.7098 0.6535 0.6776 qwen-vl-chat[4] 0.5299 0.5308 -0.0002 0.0475 0.0410 0.0569 0.9737 0.5111 shikra-7b[12] 0.4294 0.2736 -0.0848 0.0672 0.1312 0.1269 0.9148 0.5116 yi-vl[52] 0.6218 0.4688 0.4326 0.2617 0.2222 0.4086 0.8962 0.5969 InternVL3-8b[57] 0.7151 0.6062 0.2163 0.1568 0.2662 0.3601 0.9005 0.5808 kimi-vl-A3B-Instruct[42] 0.5574 0.5172 0.2548 0.1496 0.2289 0.3480 0.9159 0.5790 qwen2.5-vl-7b[6] 0.6541 0.5236 0.3598 0.1657 0.4771 0.4294 0.8196 0.5885 qwen3.0-vl-8b[5] 0.6561 0.5858 0.5469 0.1254 0.1588 0.3691 0.8757 0.5467 GPT-4o[33] 0.7501 0.6979 0.6507 0.4758 0.4776 0.6709 0.7415 0.6415 Gemini-1.5-pro[36] 0.6702 0.6842 0.4563 0.4249 0.2485 0.4657 0.7651 0.5724 Average 0.5732 0.4896 0.2223 0.1771 0.2528 0.3017 0.8915 0.5677
V-B P1: Misalignment Between Privacy Awareness and Privacy Leakage
V-B1 High-Level Evidence
We first examine aggregate results to evaluate the relationship between privacy awareness and privacy leakage. If privacy protection were primarily awareness-driven, models with higher awareness accuracy would be expected to exhibit lower leakage.
Table II shows that strong privacy awareness does not necessarily translate into reduced privacy leakage. GPT-4o [33] consistently achieves the highest level of privacy awareness across all three tasks, indicating a comparatively strong ability to recognize privacy-related content. However, this advantage does not lead to a corresponding reduction in privacy leakage, suggesting that awareness alone is insufficient to ensure effective privacy protection. Nevertheless, accurately distinguishing between privacy-relevant and privacy-irrelevant inputs remains challenging, and beyond GPT-4o [33], most models fail to show consistent alignment with human judgments in scenarios involving the propagation of private information. In contrast, phi-3-vision [2] demonstrates strong privacy preservation capabilities, ranking first in Privacy Leakage across all three tasks by effectively declining sensitive requests, while its performance in Privacy Awareness is relatively limited, especially in terms of alignment with human judgments in Privacy InfoFlow Assessment.
These observations indicate that, at the aggregate level, privacy awareness and privacy leakage are generally inconsistent.
V-B2 Fine-Grained Analysis
The Multi-PA benchmark enables a one-to-one correspondence between Privacy Question Detection and Privacy Leakage samples, allowing a direct comparison of awareness and protection behavior.